OpenStudy (anonymous):

Can any please help me Which of the following is the slope between the points (0, -10) and (-4, 0)? two fifths -two fifths -five halves five halves

OpenStudy (mertsj):

1. Subtract the y values 2. Subtract the x values 3. Put the answer to 1 over the answer to 2

OpenStudy (anonymous):

ok i did all of that @Mertsj

OpenStudy (kagıtucak):

m=y-y0/x-x0

OpenStudy (anonymous):

i did that

OpenStudy (kc_kennylau):

and what have you got

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

1/2

OpenStudy (mertsj):

What did you get when you subtracted the y values?

OpenStudy (anonymous):

-2/5 but that is not right so i am confused

OpenStudy (kc_kennylau):

y over x not x over y

OpenStudy (mertsj):

What is the y value of the first ordered pair?

OpenStudy (anonymous):

@Mertsj is this what you mean Which of the following is the slope between the points (0, -10) and (-4, 0)? two fifths -two fifths -five halves five halves

OpenStudy (mertsj):

The problem asks you to find the slope. Do you understand that?

OpenStudy (mertsj):

That is a simple yes or no question.

OpenStudy (anonymous):

yes and i got -2/5

OpenStudy (mertsj):

Do you know which number is the y value in an ordered pair?

OpenStudy (mertsj):

Another simple yes or no question.

OpenStudy (anonymous):

5

OpenStudy (mertsj):

The first ordered pair is (0,10) Which number is the y value...the 0 or the 10?

OpenStudy (anonymous):

10

OpenStudy (mertsj):

Actually the ordered pair is (0,-10). And yes, the -10 is the y value.

OpenStudy (mertsj):

The second ordered pair is (-4,0). What is the y value?

OpenStudy (anonymous):

0

OpenStudy (mertsj):

Subtract those two y values: -10-0 = ???

OpenStudy (anonymous):

it -10

OpenStudy (mertsj):

No write a fraction and put -10 on the top like this:

OpenStudy (mertsj):



OpenStudy (mertsj):

Now the x values are 0 and -4 so do that subtraction: 0-(-4)= ???

OpenStudy (anonymous):

@Mertsj are you still on here

OpenStudy (mertsj):

yes but you won't answer my questions and others need my help.

OpenStudy (anonymous):

i was wounding if you were still here

OpenStudy (anonymous):

the anwser to your question is 4 @Mertsj

OpenStudy (mertsj):

Now finish the fraction by putting that number on the bottom of the fraction.

OpenStudy (mertsj):